Alyssa was born in the (7x+18) and she will be 17 years old in the year (8x-247). in what year was Alyssa born?

Answer by stanbon(75887)

Alyssa was born in the (7x+18) and she will be 17 years old in the year (8x-247). in what year was Alyssa born?
--------------------
7x+18 + 17 = 8x-247
7x + 35 = 8x-247
x = 282
----
Ans: 7x+18 = 1992
-------------
Cheers,
Stan H.
------------